OpenStudy (anonymous):

simplify the expression "the square root of x^10 times the square root of x^4." a. the square root of x^14 b. 2 times the square root of x^10 c. 3 times the root of x^5 d. x^7

OpenStudy (king):

option d

OpenStudy (ash2326):

\[\sqrt(x^(10))*\sqrt(x^4)\] \[\sqrt(x^(10)*x^4)\] \[\sqrt(x^(14))\] \[x^7\] option d